Solve the system by graphing, if possible.
If the system is dependent or inconsistent, indicate this.
In those cases enter dependent or inconsistent. Otherwise, enter your answer as an ordered pair in the form (x, y).
{x -y =7
{3x + y =5

Answer by Nate(3500)

let us set the equations is y=mx+b form.....where m is the slope and b is the y-intercept
x -y =7
-y = -x + 7
y = x - 7
---------------------------------------------------------
3x + y =5
y = -3x + 5
the slopes are different, so they are not parallel.... that means the lines are perpendicular; if you combine the equations, you get (3,-4)